OpenStudy (anonymous):

how do i find the sum of a series where n=1 to infinite and 1/(1+(2/3)^n) ? ....

OpenStudy (shamil98):

The sum for a series of n is the following the sum of n = 1/2(first term + the last term or the term you go up to)(the term you go up to)

OpenStudy (shamil98):

for example lets say a n = n + 1 find the sum for 10 . s = 1/2(2 + 11) (10) s = 11(10) s= 121